a radio stations channel, such as 100.7 FM, is actually it's frequency in MHz, calculate the broadcast wavelength of the radio station 100.9 FM
Answer #1

100.9 FM = 100.9 MHz

100.9 MHz = 100.9 * 106 Hz

Wavelength λ = c / f =( 3*108)/(100.9 * 106 )= 2.97 meter

so broadcast wavelength = 2.97 meter